# Adversarial Disentanglement by Backpropagation with Physics-Informed Variational Autoencoder

### v1

@@ 7 Conclusions
Three synthetic case studies involving a Euler-Bernoulli beam, a damped oscillator, and a population of bridges were investigated. In these cases, a nominal model representing the partially known physics was available or built from a simulator. For each case, noisy observations of the structural response and information on domain (the environmental and operational conditions that a system is exposed to, as well as other parameters that may not be directly specified in the model of the known physics) and class (the characteristics of a structure related to the existence and extent of damage and degradation) are assumed available. It was shown that the proposed architecture promotes the learning of disentangled representations, and mitigates the issues that occur when including physics-based components in standard VAE. Furthermore, it was shown that the proposed approach is able to: (i) Preferentially utilize the known physics, resulting in an interpretable and physically meaningful posterior distribution over physics-grounded latent variables, (ii) Accurately reconstruct the structural response in the presence of domain and class influences that are not described by the known physics, and (iii) Predict the class variables associated with a structure under previously unseen conditions using noisy measurements of the structural response.
@@
The code and data (generated via the synthetic use cases) needed to replicate the results shown in this paper can be accessed through the link: https://github.com/JanKoune/DPI-VAE.
@@ Appendix A Implementation details
The encoder reduces the dimensionality of the input measurements and maps them to vectors of mean values $\bm{\mu}_{\mathrm{\phi}}(\bm{x})$, standard deviations $\bm{\sigma}_{\mathrm{\phi}}(\bm{x})$ and a lower triangular matrix $\bm{L}^{\prime}_{\mathrm{\phi}}(\bm{x})$. The lower triangular factor $\bm{L}_{\mathrm{\phi}}(\bm{x})=\bm{L}^{\prime}_{\mathrm{\phi}}(x)+\bm{\sigma}_{\mathrm{\phi}}(\bm{x})\bm{I}$ is the Cholesky decomposition factor of the covariance matrix $\bm{\Sigma}_{\mathrm{\phi}}(\bm{x})$, i.e. $\bm{\Sigma}_{\mathrm{\phi}}(\bm{x})=\bm{L_{\mathrm{\phi}}(\bm{x})L_{\mathrm{\phi}}(\bm{x})}^{T}$, such that the posterior distribution corresponding to each input is a multivariate Normal distribution. The encoder outputs the log of the standard deviations, which are then exponentiated to avoid negative values. The reparametrization trick [Kingma2022] is exploited to define the latent variables $\bm{z}$ as a deterministic transformation of a noise variable $\bm{\epsilon}\sim p(\bm{\epsilon})$. This facilitates the computation of unbiased Monte Carlo gradient estimates of the objective with respect to the variational parameters, using automatic differentiation. With the exception of the introductory examples presented in Section 2.4, the encoder is everywhere formulated as a single feed-forward NN using a shallow architecture with a single hidden layer. The input and hidden layer widths are $[d_{\mathrm{x}},128]$, where $d_{\mathrm{x}}$ is the dimensionality of the input. The output layer is composed of three heads, corresponding to the mean, standard deviation and covariance outputs. The mean and standard deviation heads have output sizes of $d_{\mathrm{z}}$, while the covariance head has an output size of $d_{\mathrm{z}}^{2}$, where $d_{\mathrm{z}}=d_{\mathrm{z_{\mathrm{x}}}}+d_{\mathrm{z_{\mathrm{c}}}}+d_{\mathrm{z_{\mathrm{y}}}}$, with $d_{\mathrm{z_{\mathrm{i}}}}$ denoting the size of the $i$’th subset of the latent space. A ReLU activation function is applied on all layers except the final output layer. For the introductory examples, three independent encoders are used instead. The hidden layer widths of each independent network are set to 128 units, and the output shapes are adjusted according to the dimensionality of the corresponding subset of the latent space. To further ensure numerical stability, the outputs of all encoder NNs are clamped within ranges of values that are reasonably expected to be encountered for the case studies investigated in this work.
@@
The decoder of the response is formulated as a feed-forward NN with a single, 128-unit-wide hidden layer and a ReLU nonlinearity at the output of the hidden layer. The size of the input is $d_{\mathrm{z}_{\mathrm{c}}}+d_{\mathrm{z}_{\mathrm{y}}}$ and the output size is $d_{\mathrm{x}}$. A gradient reversal layer is placed at the input of this network. For the structural response prediction, the standard deviation $\sigma_{\mathrm{x}}$ is included in the vector $\bm{\theta}$ and jointly optimized with the NN hyperparameters. The auxiliary networks are similarly formulated with a single hidden layer with a width of $128$ units and a ReLU nonlinearity between the input and the hidden layer. The auxiliary decoders are composed of two prediction heads, responsible for the mean prediction and standard deviation respectively. The input and output shapes are $d_{\mathrm{z}_{i}}$ and $2\cdot d_{i}$, where $i$ denotes the corresponding domain or class modality.
@@
The conditional prior distribution networks share the same formulation as the encoder networks, but use $128$ units in the hidden layer. The input and output shapes are also adjusted to $d_{\mathrm{i}}$ and $d_{\mathrm{z}_{i}}$ respectively, where the subscript $i\in\{x,c,y\}$ denotes the corresponding modality and subset of the latent space.
@@
Optimization is carried out using the Adam algorithm [Kingma2017] with minibatch gradient estimation [Kingma2022]. The model is trained for up to $20,000$ epochs with a batch size equal to the size of the full training dataset $N_{\mathrm{train}}=1024$, which is possible due to the relatively small size of the datasets used in the case studies presented in this work. Early stopping is implemented by monitoring the value of the ELBO, evaluated on a held-out validation set with size $N_{\mathrm{val}}=512$. The training is terminated if no improvement of the ELBO is observed over $2,000$ epochs. For the experiments described in this work, convergence was typically achieved within $5,000$ to $10,000$ epochs. Gradient and objective estimates are obtained using $8$ Monte Carlo samples during training and $512$ samples during testing, although in practice the training was found to be insensitive to the number of samples, likely due to the large batch size. All learning rates are set to $0.001$, except for the learning rate of the standard deviation parameter for the response $\sigma_{\mathrm{x}}$, which is set to $0.005$.
@@
Figures illustrating the traversal of the latent space and the space of reconstructions are provided for each case study. Samples from the latent space and reconstructions are obtained as follows: Five linearly spaced values between the $5^{\mathrm{th}}$ and $95^{\mathrm{th}}$ percentile of the ground truth distribution are computed for each generative factor in turn, while the remaining generative factors are fixed to a constant value. For each combination of generative factors, $1000$ realizations of response measurements are generated using the procedure described in Section 5. The model is evaluated on the response measurements, and a single sample is drawn from the approximate posterior distribution for each response measurement. The decoder is then evaluated on each sample from the posterior, yielding deterministic predictions $\hat{\bm{x}}_{\mathrm{p}}$ and $\hat{\bm{x}}_{\mathrm{d}}$ from the physics-based and data-driven components respectively. The combined prediction is sampled from $\mathcal{N}(\hat{\bm{x}}_{\mathrm{p}}+\hat{\bm{x}}_{\mathrm{d}},\sigma^{2}_{\mathrm{x}}\bm{I})$. The visualizations of the latent space and reconstructions therefore also include the randomness in the data generating process, in addition to the randomness in the approximate posterior distribution and decoder.

### v2

@@ 7 Conclusions
Three synthetic case studies involving a beam, an oscillator, and a population of bridges were investigated. In these cases, a nominal model representing the partially known physics was available or built from a simulator. For each case, noisy observations of the structural response and information on domain (the environmental and operational conditions that a system is exposed to) and class (the characteristics of a structure related to the existence and extent of damage and degradation) are assumed available. It was shown that the proposed architecture promotes the learning of disentangled representations, and mitigates the issues that occur when including physics-based components in standard VAE. Furthermore, it was shown that the proposed approach is able to: (i) Preferentially utilize the known physics, resulting in an interpretable and physically meaningful posterior distribution over physics-grounded latent variables, (ii) Accurately reconstruct the structural response in the presence of domain and class influences that are not described by the known physics, and (iii) Predict the class variables associated with a structure under previously unseen conditions using noisy measurements of the structural response.
@@
The code and data (generated via the synthetic use cases) needed to replicate the results shown in this paper can be accessed through the link: https://doi.org/10.5281/zenodo.15813028.
@@ Appendix A Implementation details
The encoder reduces the dimensionality of the input measurements and maps them to vectors of mean values $\bm{\mu}_{\mathrm{\phi}}(\bm{x})$, standard deviations $\bm{\sigma}_{\mathrm{\phi}}(\bm{x})$ and a lower triangular matrix $\bm{L}^{\prime}_{\mathrm{\phi}}(\bm{x})$. The lower triangular factor $\bm{L}_{\mathrm{\phi}}(\bm{x})=\bm{L}^{\prime}_{\mathrm{\phi}}(x)+\bm{\sigma}_{\mathrm{\phi}}(\bm{x})\bm{I}$ is the Cholesky decomposition factor of the covariance matrix $\bm{\Sigma}_{\mathrm{\phi}}(\bm{x})$, i.e. $\bm{\Sigma}_{\mathrm{\phi}}(\bm{x})=\bm{L_{\mathrm{\phi}}(\bm{x})L_{\mathrm{\phi}}(\bm{x})}^{T}$, such that the posterior distribution corresponding to each input is a multivariate Normal distribution. The encoder outputs the log of the standard deviations, which are then exponentiated to avoid negative values. The reparametrization trick [Kingma2022] is exploited to define the latent variables $\bm{z}$ as a deterministic transformation of a noise variable $\bm{\epsilon}\sim p(\bm{\epsilon})$. This facilitates the computation of unbiased Monte Carlo gradient estimates of the objective with respect to the variational parameters, using automatic differentiation. With the exception of the introductory examples presented in Section 2.4, the encoder is everywhere formulated as a single feed-forward NN using a shallow architecture with a single hidden layer. The input and hidden layer widths are $[d_{\mathrm{x}},128]$, where $d_{\mathrm{x}}$ is the dimensionality of the input. The output layer is composed of three heads, corresponding to the mean, standard deviation and covariance outputs. The mean and standard deviation heads have output sizes of $d_{\mathrm{z}}$, while the covariance head has an output size of $d_{\mathrm{z}}^{2}$, where $d_{\mathrm{z}}=d_{\mathrm{z_{\mathrm{x}}}}+d_{\mathrm{z_{\mathrm{c}}}}+d_{\mathrm{z_{\mathrm{y}}}}$, with $d_{\mathrm{z_{\mathrm{i}}}}$ denoting the size of the $i$’th subset of the latent space. A ReLU activation function is applied on all layers except the final output layer. For the introductory examples, an independent encoder network is used for each subset of the latent variables. The hidden layer widths of each independent network are set to 64 units, and the output shapes are adjusted according to the dimensionality of the corresponding subset of the latent variables. To further ensure numerical stability, the outputs of all encoder NNs are clamped within ranges of values that are expected to be encountered for the case studies investigated in this work.
@@
The decoder of the response is formulated as a feed-forward NN with a single, 128-unit-wide hidden layer and a ReLU nonlinearity at the output of the hidden layer. The size of the input is $d_{\mathrm{z}_{\mathrm{c}}}+d_{\mathrm{z}_{\mathrm{y}}}$ and the output size is $d_{\mathrm{x}}$. A gradient reversal layer is placed at the input of this network. For the structural response prediction, the standard deviation $\sigma_{\mathrm{x}}$ is included in the vector $\bm{\theta}$ and jointly optimized with the NN hyperparameters. The auxiliary networks are formulated with a single hidden layer with a width of $64$ units and a ReLU nonlinearity between the input and the hidden layer. The auxiliary decoders are composed of two prediction heads, responsible for the mean prediction and standard deviation respectively. The input and output shapes are $d_{\mathrm{z}_{i}}$ and $2\cdot d_{i}$, where $i$ denotes the corresponding domain or class modality.
@@
The conditional prior distributions are formulated as factorized Gaussian distributions. The corresponding neural networks use a single hidden layer with a width of $64$ units. The input and output shapes are also adjusted to $d_{\mathrm{i}}$ and $d_{\mathrm{z}_{i}}$ respectively, where the subscript $i\in\{x,c,y\}$ denotes the corresponding modality and subset of the latent space.
@@
Optimization is carried out using the Adam algorithm [Kingma2017] with minibatch gradient estimation [Kingma2022]. The model is trained for up to $20,000$ iterations with a batch size of $64$. Early stopping is implemented by monitoring the value of the ELBO, evaluated on a held-out validation set with size $N_{\mathrm{val}}=512$. The training is terminated if no improvement of the ELBO is observed over $2,000$ iterations. Gradient and objective estimates are obtained using $16$ Monte Carlo samples during training, $64$ during validation, and $512$ during evaluation, although in practice the training was found to be insensitive to the number of samples. All learning rates are set to $0.001$, except for the learning rate of the standard deviation parameter for the response $\sigma_{\mathrm{x}}$, which is set to $0.005$.
@@
Figures illustrating the traversal of the latent space and the space of reconstructions are provided for each case study. Samples from the latent space and reconstructions are obtained as follows: Five linearly spaced values between the $1^{\mathrm{st}}$ and $99^{\mathrm{th}}$ percentile of the ground truth distribution are computed for each generative factor in turn, while the remaining generative factors are fixed to a constant value. For each combination of generative factors, $1000$ realizations of response measurements are generated using the procedure described in Section 5. The model is evaluated on the response measurements, and a single sample is drawn from the approximate posterior distribution for each response measurement. The decoder is then evaluated on each sample from the posterior, yielding deterministic predictions $\hat{\bm{x}}_{\mathrm{p}}$ and $\hat{\bm{x}}_{\mathrm{d}}$ from the physics-based and data-driven components respectively. The combined prediction is sampled from $\mathcal{N}(\hat{\bm{x}}_{\mathrm{p}}+\hat{\bm{x}}_{\mathrm{d}},\sigma^{2}_{\mathrm{x}}\bm{I})$. The visualizations of the latent space and reconstructions therefore also include the randomness in the data generating process, in addition to the randomness in the approximate posterior distribution and decoder.